# Decoding Generalization from Memorization in Deep Neural Networks

### v1

@@ Abstract
Overparameterized Deep Neural Networks that generalize well have been key to the dramatic success of Deep Learning in recent years. The reasons for their remarkable ability to generalize are not well understood yet. It has also been known that deep networks possess the ability to memorize training data, as evidenced by perfect or high training accuracies on models trained with corrupted data that have class labels shuffled to varying degrees. Concomitantly, such models are known to generalize poorly, i.e. they suffer from poor test accuracies, due to which it is thought that the act of memorizing substantially degrades the ability to generalize. It has, however, been unclear why the poor generalization that accompanies such memorization, comes about. One possibility is that in the process of training with corrupted data, the layers of the network irretrievably re-organize their representations in a manner that makes generalization difficult. The other possibility is that the network retains significant ability to generalize, but the trained network somehow “chooses” to readout in a manner that is detrimental to generalization. Here, we provide evidence for the latter possibility by demonstrating, empirically, that such models possess information in their representations for substantially improved generalization, even in the face of memorization. Furthermore, such generalization abilities can be easily decoded from the internals of the trained model, and we build a technique to do so from the outputs of specific layers of the network. We demonstrate results on multiple models trained with a number of standard datasets.
@@ 1 Introduction
Prior to the advent of Deep Learning, the conventional wisdom for long111von Neumann famously said, “With four parameters I can fit an elephant, and with five I can make him wiggle his trunk.” (Dyson et al., 2004), was that in building a predictive model, the model should have as few parameters as possible and this number should certainly be less than the number of training samples that one was fitting. The dogma was that, otherwise, the model would exactly fit the training points, but invariably generalize poorly to unseen data, i.e. overfit. This intuition was also largely borne out by the models of the day. Modern Deep Learning, however, has gone on to show the opposite, namely that overparameterized models not only don’t necessarily overfit, but that they can generalize remarkably well to unseen data. However, over a decade later, we still do not satisfactorily understand why this is so. Interestingly, it has been shown (Zhang et al., 2017; 2021) that when one shuffles class labels of data points from standard training datasets to varying degrees, deep networks can still have high/perfect training accuracy on such corrupted training data; however, this appears to typically be accompanied by poor performance on unseen test data (that have true labels). This phenomenon has been called memorization, since it is thought that the model rote-learned the training data without acquiring the ability to generalize to unseen examples. It has been known (Arpit et al., 2017) that early on in training, such models start off by having better generalization ability; however generalization worsens as training accuracy increases across epochs of training. In such trained models, there have been efforts to spatially localize the origin of memorization. While certain studies (Cohen et al., 2018; Stephenson et al., 2021) have suggested that memorization occurs in the latter layers of the network, more recent work (Maini et al., 2023) suggests that memorization occurs in all layers of the network, and is largely attributable to a few units. Indeed, there have been suggestions that for networks trained on real-world data (that isn’t deliberately corrupted), memorization could play a critical role in their extraordinary performance (Feldman, 2020). While there is early work studying memorization in this setting (Feldman & Zhang, 2020), there are methodological issues that have slowed progress. Progress on understanding memorization in corrupted models could enable a better understanding of memorization and generalization in deep networks trained on real-world data.
@@
An open question arising in this context is about the detailed mechanisms that lead to poor generalization in models trained with shuffled labels. A natural hypothesis governing such mechanisms, stated informally, is that, during training, the network organizes class-conditional representations, in a manner suited to doing well on the (corrupted) training data. Since this data is significantly noisy, on being given unseen data with true labels, it fundamentally lacks the ability to have good prediction performance, leading to poor generalization. An alternative hypothesis is that layerwise representations in the network retain the ability to generalize easily, but that the network somehow chooses to readout in favor of high training accuracy in a manner that incidentally causes poor generalization performance. A corollary to this alternative hypothesis is that one ought to be able to construct a decoder for the outputs of the network’s layers that does well on the corrupted training data, while also having good generalization, i.e. high test set accuracy.
@@
Here, surprisingly, we show evidence for this alternative hypothesis. In particular, we study the organization of subspaces of class-conditioned training data on layerwise outputs of a number of deep networks. We estimate these subspaces using Principal Components Analysis (PCA). In order to remain agnostic to the information decoded by subsequent layers, we build a simple classifier that leverages the geometry of the layer output of an incoming datapoint, relative to these class-conditioned subspaces. Specifically, we measure the angle between this output vector and its projection on each of these class-conditioned subspaces and the classifier predicts this datapoint’s class to be the class whose subspace has the minimum such angle. A useful consequence of this formulation is the following. The existence of class-conditioned subspaces estimated from training data on which the aforementioned minimum angle subspace classifier has good test accuracy implies that the deep network can, in principle, generalize well.
1.1 Main Contributions
  • For models trained using standard methods & datasets with training data corrupted by label noise, while the model has poor test accuracy, we can build a simple classifier with dramatically better test accuracy that uses only the model’s hidden layer outputs obtained for the (corrupted) training set.
@@ 1.1 Main Contributions
  • For the aforementioned models, if the true training class labels are known post hoc, i.e. after the model is trained, we can build a simple classifier, with significantly better generalization performance than in (1). This is true, in many cases, even for models where training class labels are shuffled with equal probability. This demonstrates that the layers of the network maintain representations in a manner that is amenable to straightforward generalization to a degree not previously recognized.
@@
  • On the other hand, we asked if a model trained on the true training labels similarly retained the capability to memorize easily. Adapting our technique to this setting, we find that in a few cases, we can extract a high degree of memorization. The same classifier sometimes exhibits high test accuracy (on the true test labels), which further supports the idea that generalization can co-exist with memorization.
@@ 2 Related work
The idea of probing intermediate layers of Deep Networks isn’t new. For example, (Montavon et al., 2011; Alain & Bengio, 2018) do so by using kernel PCA & linear classifiers respectively. However, this approach has not been used to investigate memorization. Indeed, (Alain & Bengio, 2018) explicitly avoid examining memorized networks from (Zhang et al., 2017) because they thought such probes would inevitably overfit. Our results are therefore especially surprising in this context.
There is evidence that DNN’s learn simple patterns first, before memorizing (Arpit et al., 2017), & DNNs learn lower frequencies first (Belrose et al., 2024). (Stephenson et al., 2021) study memorized models, concluding that memorization happens in later layers, since rewinding early layer weights to their early stopping values recovers some generalization, but rewinding later layer weights doesn’t. On the contrary, our results suggest that later layers in most models investigated retain significant ability to generalize, & we demonstrate this without modifying the weights of the trained network.
@@
There is an important line of theoretical work in deep linear models (Saxe et al., 2013) where the question of generalization has been studied. In this context, (Lampinen & Ganguli, 2018) offer a theoretical explanation for the phenomenon of memorization in networks trained with noisy labels.
@@
Experiments towards understanding training dynamics across layers using different Canonical Correlation Analysis have be explored (Raghu et al., 2017) and in various generalized and memorized networks is analyzed (Morcos et al., 2018). Centered Kernel Alignment in different random initializations by (Kornblith et al., 2019) and network similarity between model trained with same data and different initialization is examined by (Wang et al., 2018). Also experiments related to using measures of the representational geometry towards understanding dynamics of layerwise outputs (Chung et al., 2016; Cohen et al., 2020) and different measures such as curvature (Hénaff et al., 2019) Dimensionality which express the structures within the representations (Sussillo & Abbott, 2009; Farrell et al., 2019; Gao & Ganguli, 2015; Litwin-Kumar et al., 2017; Bakry et al., 2015; Cayco-Gajic & Silver, 2019; Yosinski et al., 2014; Stringer et al., 2019; Yosinski et al., 2014) have also been explored.
@@
To deal with label noise, many heuristics have been explored (Khetan et al., 2017; Scott et al., 2013; Reed et al., 2014; Zhang & Sabuncu, 2018; Malach & Shalev-Shwartz, 2017) & for classification task see (Frénay et al., 2014; Ren et al., 2018; Menon et al., 2018; Shen & Sanghavi, 2019). For over parameterized models, (Li et al., 2020) shows that the memorized network weights are far away from the initial random state in order for them to overfit the noisy labels. (Stephenson & Lee, 2021) propose a theoretical model for epochwise double descent that suggests that for small-sized models, moderate amounts of noise can cause generalization error to dip later on in training.
3 Methodology
Using the organization of subspaces of class-conditioned training data on layerwise outputs of deep networks, we build a Minimum Angle Subspace Classifier (MASC) with the following steps:
Creation of subspaces: For a specific layer, we estimate subspaces for each class. The class-conditioned training data subspaces on layerwise outputs of deep networks are computed using PCA. If the empirical mean of the class-conditioned data isn’t zero, PCA in effect, will provide us an affine space, i.e. a linear space that doesn’t pass via the origin. However, we have determined subspaces – which are linear spaces passing through the origin – here rather than affine spaces. In order to do so, we add the negative of each sample to the dataset so it is guaranteed to have empirical mean be zero, before running PCA. This created dataset is sent to the PCA algorithm to calculate PCA components for a certain percentage of variance explained in the dataset. The span of these PCA components is the subspace $S$. We illustrate the process for a Multi-layer Perceptron (MLP) model in Figure 4 in the Appendix.
Projection of the data point: Layer output of an incoming data point is projected onto these class-specific subspaces.
Label assignment using minimum angle: For every data point, the angle between the original data point and projected data point for each class is calculated. The Minimum Angle Subspace Classifier (MASC) assigns to the datapoint, the label of the subspace having the minimum angle with the original data point.
While the subspaces are estimated using the training data alone, accuracy of the Minimum Angle Subspace Classifier is determined for the training data and the testing data separately. This process is followed for all the layers in the network independently. MASC is using labels of the dataset while creating the class-specific subspaces. For experiments in Section 4, MASC uses corrupted training labels whereas in Section 5, MASC uses true training labels to create class-specific subspaces. See Appendix A.7 for MASC algorithm. We have used 99% as the percentage of variance explained, unless otherwise mentioned.
@@ 3.1 Experimental Setup
We have used multiple models and datasets, namely Multi-layer Perceptron (MLP) trained on MNIST (Deng, 2012) and CIFAR-10 (Krizhevsky, 2009) datasets, Convolutional Neural Networks (CNN) 222 The CNN models were built along the lines of (Tran et al., 2022). trained on MNIST, Fashion-MNIST (Xiao et al., 2017), and CIFAR-10 and AlexNet (Krizhevsky et al., 2012) trained on CIFAR-100 (Krizhevsky, 2009) and Tiny ImageNet (Moustafa, 2017). We have trained these models with training data having true labels (“generalized models”) as well as separately using training data with labels shuffled to varing degrees (“memorized models”) (Zhang et al., 2021).
For memorized models, when we say we train it with corruption degree $p$, we mean that with probability $p$, we attempt changing the label for a training datapoint. Changing the labels happens uniformly at random. Note that this may result in the label remaining the same; therefore the expected fraction of datapoints whose labels changed are $p-p/c$ where $c$ is the number of classes. So, this would mean that for corruption degrees of 20% , 40%, 60%, 80%, 100% the expected percentage of training datapoints with changed labels is 18%, 36%, 54%, 72%, 90% respectively, when $c=10$. We have run experiments for values of $p$ being 0% (generalized model), 20% , 40%, 60%, 80%, 100% (memorized models).
@@
A summary of the models and datasets with training set size and number of parameters is in Table 1 in the Appendix. The average training and testing accuracies of all the models over three runs are shown in Table 2 and 3 in section A.3. More details of these models, hyperparameters & training are available in Section A.2. Following standard practice in probing memorized models (e.g. (Stephenson et al., 2021)), we do not use explicit regularizers such as Dropout or batchnorm, or early stopping, unless otherwise mentioned, as a result of which our baseline test accuracy numbers are often much lower than what is usually found with standard training of these models. All the models are trained to either reach very high training accuracies (i.e. $99\%-100\%$) or trained until 500 epochs. Some models did not result in such high accuracies, in which case, results have been shown on the model obtained at epoch 500. We trained 3 instances of each model and results displayed are averaged over these instances with the shaded region indicating the range of results also indicated in the plots.

### v2

@@ Abstract
Overparameterized deep networks that generalize well have been key to the dramatic success of deep learning in recent years. The reasons for their remarkable ability to generalize are not well understood yet. When class labels in the training set are shuffled to varying degrees, it is known that deep networks can still reach perfect training accuracy at the detriment of generalization to true labels – a phenomenon that has been called memorization. It has, however, been unclear why the poor generalization to true labels that accompanies such memorization, comes about. One possibility is that during training, all layers of the network irretrievably re-organize their representations in a manner that makes generalization to true labels difficult. The other possibility is that one or more layers of the trained network retain significantly more latent ability to generalize to true labels, but the network somehow “chooses” to readout in a manner that is detrimental to generalization to true labels. Here, we provide evidence for the latter possibility by demonstrating, empirically, that such models possess information in their representations for substantially-improved generalization to true labels. Furthermore, such abilities can be easily decoded from the internals of the trained model, and we build a technique to do so. We demonstrate results on multiple models trained with standard datasets. Our code is available at: https://github.com/simranketha/MASC_DNN.
@@ 1 Introduction
Prior to the advent of deep learning, the conventional wisdom for long111von Neumann famously said, “With four parameters I can fit an elephant, and with five I can make him wiggle his trunk.” (Dyson and others, 2004), was that in building a predictive model, the model should have as few parameters as possible and this number should certainly be less than the number of training samples that one was fitting. The dogma was that, otherwise, the model would exactly fit the training points, but invariably generalize poorly to unseen data, i.e. overfit. This intuition was also largely borne out by the models of the day. Modern deep learning, however, has gone on to show the opposite, namely that overparameterized models not only don’t necessarily overfit, but that they can generalize remarkably well to unseen data. However, over a decade later, we still do not satisfactorily understand why this is so. Interestingly, it has been shown (Zhang et al., 2017; 2021) that when one randomly shuffles class labels of data points from standard training datasets to varying degrees, deep networks can still have high/perfect training accuracy when trained on such corrupted training data; however, this appears to typically be accompanied by poor performance on unseen test data (that have true labels). This phenomenon has been called memorization [^2] We direct the reader to Section 3.1 for a formal description of our setting., since it is thought that the model rote-learned the training data without acquiring the ability to generalize to true labels. It has been suggested that progress on understanding memorization could enable a better understanding of generalization to true labels in deep networks trained on real-world data (Zhang et al., 2017; 2021) and indeed that a detailed understanding of mechanisms of generalization to true labels should also be able to explain the phenomenon of memorization.
@@
An open question arising in this context is about the detailed mechanisms that lead to poor generalization to true labels in models trained with shuffled labels, i.e. models that memorize. A natural hypothesis governing such mechanisms, stated informally, is that, during training, the network organizes its internal representations in all layers, in a manner suited to doing well on the (corrupted) training data. Since this data is significantly noisy, on being given unseen data with true labels, it fundamentally lacks the ability to have good prediction performance, leading to poor generalization to true labels. An alternative hypothesis is that layerwise representations on a subset [^3] It is indeed possible that certain layers of the network have representations that are significantly more generalizable to true labels than others and these layers may be early or later layers. of the layers in the network retain significantly more ability to generalize to true labels, than the model, but that the network somehow chooses to readout in favor of high training accuracy in a manner that incidentally causes poor generalization to true labels. A consequence of this alternative hypothesis is that one ought to be able to construct a decoder (i.e. a probe) for the outputs of such layers that has better generalization performance on true labels.
@@
Here, surprisingly, we show evidence for this alternative hypothesis. In particular, we study the organization of subspaces of class-conditioned training data on layerwise outputs, in deep networks. We estimate these subspaces using Principal Components Analysis (PCA). In order to remain oblivious to the information decoded by subsequent layers, we build a simple probe that leverages the geometry of the present layer’s output of an incoming datapoint, relative to these class-conditioned subspaces. Specifically, we measure the angle between this output vector and its projection on each of these class-conditioned subspaces and the probe predicts the datapoint’s class to be the class whose subspace has the minimum such angle. We call this probe the Minimum Angle Subspace Classifier (MASC). Notably, unlike probes used conventionally (e.g. in (Alain and Bengio, 2018)) whose parameters are determined by iteratively minimizing a crossentropy loss, the parameters of MASC are directly determined from the subspace geometry of the training data. A schematic illustrating the geometry of MASC is presented in Figure 2.
Figure 1: A schematic of the memorization setting used in our work and the application of the MASC classifier in it. a) Illustration of a standard training dataset. b) A corrupted training dataset is created by changing the labels of the standard training dataset with a specific probability (due to which a few of the colors are changed, representing the changed labels). Changing the labels happens uniformly at random for the whole dataset. c) A deep network is trained with this corrupted training dataset to achieve $\sim$100% training accuracy, which is usually accompanied by poor test accuracy (as measured on true labels from the test set). We have shown that the Minimum Angle Subspace Classifier (MASC) – our technique – which uses the internals of the deep network, tends to have significantly better generalization to true labels (test accuracy) than the deep network itself.
We train a number of deep networks with standard datasets in the memorization setting. Here, a randomly-chosen fraction of training data points have their labels changed to a randomly-chosen label from the available labels in the dataset. We do so for differing fractions of the training dataset and – consistent with previous work (Zhang et al., 2017; 2021; Arpit et al., 2017) – see that training with such corrupted training datasets causes correspondingly poor test accuracies in the model. However, MASC – which uses the internals of the network to predict the class label – tends to do significantly better [^4] We find that a few other probes do comparably well. than the model on the test set. A schematic illustration of the memorization setting with MASC is shown in Figure 1.
We outline a more detailed summary of our main contributions below.
  1. 1.
    For models trained with standard methods & datasets with training data corrupted by label noise to varying degrees, we demonstrate (with one exception) that MASC applied on at least one layer, when using subspaces corresponding to such corrupted training data, has significantly better test accuracy than the model. For example, MASC outperforms the model test accuracy by upto 159.93%, 189.00%, 64.86% and 119.16% on MLP-MNIST, CNN-Fashion-MNIST, AlexNet-CIFAR-100 and ResNet-18-CIFAR-10 respectively. A more detailed account of these numbers is in Table 1.
  2. 2.
    For the models discussed above, we perform a comparison study evaluating five probes namely Logistic regression , K-Nearest Neighbor (KNN), Linear Discriminant Analysis(LDA), Quadratic Discriminant Analysis (QDA), and Nearest Class Mean (NCM) over the layers of multiple deep networks under varying corruption degrees. While many of these probes exhibit similar overall accuracies and show no consistent trends across corruption degrees, their computational costs differ considerably, which we have empirically analyzed.
  3. 3.
@@
    For the aforementioned models, if the true training class labels are known post hoc, i.e. after the model is trained, we can build MASC using subspaces corresponding to true class labels. These MASC classifiers usually have better generalization to true labels than in (1). For example, MASC using true labels outperforms the model by upto 198.43%, 212.42%, 337.51% and 228.64% on MLP-MNIST, CNN-Fashion-MNIST , AlexNet-CIFAR-100 and ResNet-18-CIFAR-10 respectively. A more detailed account of these numbers is in Table 7. This demonstrates that the layers of the memorized network maintain representations in a manner that is amenable to straightforward generalization to true labels to a degree not previously recognized.
  4. 4.
@@
    Conversely, we asked if a model trained on true training labels similarly retained internal representations that have the capability to memorize easily, as manifested by MASC. Adapting our technique to this setting, we create corrupted training sets which we use to build MASC. In this setting, we find that we can extract a high degree of memorization, in some cases. The results are presented in Section H.
  5. 5.
    Finally, leveraging the MASC classifiers built in (1) and (3), we ask, if we can retrain the memorized model for a few epochs to achieve better model generalization to true labels. We find that indeed, in many cases, there is an improvement in the generalization to true labels of the model.
@@ 2 Related work
The idea of probing intermediate layers of deep networks isn’t new. For example, kernel-PCA (Montavon et al., 2011) with RBF kernels has been used to analyze layerwise evolution of representations of deep networks. In that work, they quantify the quality of layerwise representations and find that the last layers of the network tend to have representations that are more simple and accurate than previous layers. Likewise, linear classifier probes (Alain and Bengio, 2018) have been used to study the roles and dynamics of intermediate layers in deep networks. There, they show that the degree of linear separability increases over the layers of the network. However, they explicitly avoid examining memorized networks (Zhang et al., 2017) because they thought such probes would inevitably overfit. Our results are therefore especially surprising in this context, because we demonstrate, on the contrary, that intermediate representations, in fact, tend to resist overfitting, to a degree not previously recognized.
The authors in (Zhang et al., 2017) investigate the question of whether deep networks can perfectly learn noise by training on randomly relabeled data, showing that although such models achieve near-perfect training accuracy, they exhibit poor generalization to the true labels. Subsequent work by (Arpit et al., 2017) shows that early on in training, memorized networks (Zhang et al., 2017) start off by having better generalization to true labels; however generalization to true labels worsens as training accuracy increases across epochs of training. Despite these observations, there have been limited efforts to analyze the internal representations of models in the memorization regime (i.e. one where the labels of a subset of training data points are shuffled randomly) using probing methods. It was assumed that the cause of poor generalization to true labels in memorization settings directly manifests from poor representations. For example, in a probe-based analysis (Alain and Bengio, 2018) argue that probe measurements would be “entirely meaningless” in memorization settings. They say
“It was recently demonstrated that very large models can fit random labels on ImageNet (Zhang et al., 2016). This is a situation that we want to avoid because the probe measurements would be entirely meaningless in that situation.”
This view aligns with our null hypothesis that, during training, all layers of the network irretrievably re-organize their representations in a manner that makes generalization to true labels difficult. However, the alternative hypothesis that we articulate here has not been explicitly examined in past work, presumably because it was not considered plausible: that one or more layers of the trained network retain significantly more latent ability to generalize to true labels, but the network somehow “chooses” to readout in a manner that is detrimental to generalization to true labels.
There have been efforts to build training algorithms that are designed to extract better generalization to true labels in the case where the data is known to be noisy (Jiang et al., 2018; Han et al., 2018; Liu et al., 2020). Stephenson et al (Stephenson et al., 2021) investigate memorized models, suggesting that memorization predominantly occurs in the later layers. This is based, in part, on the observation that rewinding early-layer weights to their early-stopping values can recover generalization to true labels, whereas rewinding later-layer weights does not yield the same effect. In general, the thinking in the field has been that while there is an initial peak in generalization to true labels, it is lost during further training, although one can mitigate some of this loss by modifying training (Jiang et al., 2018) or by rewinding a subset of weights to their early values. On the contrary, our results suggest that layerwise outputs of deep networks retain significant ability to generalize after training and we demonstrate that this generalization to true labels can be extracted without modifying the weights of the trained network that are obtained via standard training methods.
@@
An important line of theoretical research on deep linear models has explored the question of generalization to true labels (Saxe et al., 2013). Here, a theoretical explanation for the phenomenon of memorization in networks trained with noisy labels has been proposed (Lampinen and Ganguli, 2018).
@@
Studies have investigated training dynamics across layers using various forms of Canonical Correlation Analysis (Raghu et al., 2017), including analyses in both generalized and memorized networks (Morcos et al., 2018). Centered Kernel Alignment has been employed to examine the effects of different random initializations (Kornblith et al., 2019), as well as to study network similarity between models trained on the same data with different initializations (Kornblith et al., 2019). Additionally, experiments have explored the use of representational geometry measures to understand the dynamics of layerwise outputs (Chung et al., 2016; Cohen et al., 2020), along with other structural measures such as curvature dimensionality (Hénaff et al., 2019), which aim to capture underlying properties of learned representations (Sussillo and Abbott, 2009; Farrell et al., 2019; Gao and Ganguli, 2015; Litwin-Kumar et al., 2017; Bakry et al., 2015; Cayco-Gajic and Silver, 2019; Yosinski et al., 2014; Stringer et al., 2019).
@@
To address label noise, various heuristic approaches have been proposed (Khetan et al., 2017; Scott et al., 2013; Reed et al., 2014; Zhang and Sabuncu, 2018; Malach and Shalev-Shwartz, 2017), particularly in the context of classification tasks (Frénay et al., 2014; Ren et al., 2018; Menon et al., 2018; Shen and Sanghavi, 2019). In the case of overparameterized models, Li et al (Li et al., 2020) demonstrate that memorization requires the network weights to deviate significantly from their initial random state in order to overfit noisy labels. Additionally, in a theoretical model of epochwise double descent (Stephenson and Lee, 2021), it has been suggested that for smaller models, moderate levels of label noise can lead to a reduction in generalization error at later stages of training.
3 Methods
3.1 Preliminaries
In this subsection, we state precisely the setting that is treated in this paper.
We study a classification task defined over a data distribution $D$, with an i.i.d. training dataset $T$ drawn from $D$. For a given corruption degree pp [^5] When we say the training dataset has corruption degree $p$, we mean that with probability $p$, we attempt changing the label for each training datapoint. Changing the labels happens uniformly at random to any of the class labels. Note that this may result in the label remaining the same; therefore the expected fraction of datapoints whose labels changed are pp/Kp-p/K, where $K$ is the number of class labels. So, e.g. for a dataset with $10$ classes, this would mean that for corruption degrees of 20%, 40%, 60%, 80% and 100%, the expected percentage of training datapoints with changed labels is 18%, 36%, 54%, 72% and 90% respectively. We have run experiments for values of $p$ being 0% (generalized model), and memorized models with $p$ being 20%, 40%, 60%, 80% and 100% ., we generate a modified training set $\hat{T}_{p}$ by randomly relabeling a certain expected fraction of the training samples uniformly at random. The resulting label corrupted training data distribution is denoted by $\hat{D}_{p}$.
A series of prior studies (e.g., Zhang et al. (2017); Arpit et al. (2017)) have demonstrated that deep networks can successfully fit such corrupted datasets $\hat{T}_{p}$, even when a substantial fraction of the labels are randomized. These findings highlight the remarkable expressive power of the hypothesis class $H$ associated with contemporary deep models. In particular, they show that $H$ is sufficiently rich to learn training data arising from highly perturbed label distributions such as $D_{p}$, thereby enabling the model to achieve near perfect training accuracy despite the presence of significant label noise.
As the corruption degree $p$ increases, the distribution of the corrupted data $\hat{D}_{p}$ diverges progressively from the true distribution $D$. Consequently, one would expect that any network $h$ trained to fit samples $\hat{T}_{p}$ drawn from $\hat{D}_{p}$ becomes increasingly misaligned with the task defined by $D$, and therefore performs poorly on test dataset $T^{\prime}$ drawn from distribution $D$. This phenomenon has been called [^6] This phenomenon could also be viewed as learning under label noise. However, given the usage of the term memorization to refer to this phenomenon in past work, we choose to continue to do so here. memorization in past work (e.g. (Zhang et al., 2017; 2021; Arpit et al., 2017)).
For a network trained on $\hat{T}_{p}$ drawn from $\hat{D}_{p}$, generalization would, usually, refer to network’s performance on $\hat{D}_{p}$ after learning from $\hat{T}_{p}$. In this work, however, we examine whether a network trained on $\hat{T}_{p}$ (using the network’s internal representations) can perform well on test data drawn i.i.d. from the true distribution $D$. Accordingly, we define good generalization in terms of high performance on a test set $T^{\prime}$ sampled from $D$. Throughout the rest of the paper, we refer to this notion simply as generalization to true labels.
@@ 3.2 Experimental setup
We have used multiple models and datasets, namely Multi-layer Perceptron (MLP) trained on MNIST (Deng, 2012) and CIFAR-10 (Krizhevsky, 2009) datasets, Convolutional Neural Networks (CNN)  [^7] The CNN models were built along the lines of (Tran et al., 2022). trained on MNIST, Fashion-MNIST (Xiao et al., 2017), and CIFAR-10, AlexNet (Krizhevsky et al., 2012) trained on CIFAR-100 (Krizhevsky, 2009) and Tiny ImageNet (Moustafa, 2017) and ResNet-18(He et al., 2016) trained on CIFAR-10. We have trained these models with training data having true labels (“generalized models”) as well as separately using training data with labels shuffled to varing degrees (“memorized models”) (Zhang et al., 2017; 2021).
A summary of the models, datasets, training set sizes, and number of parameters is provided in Table 2. Tables 3 and 4 report the average training and test accuracies of all models over three runs. Additional details on the models, hyperparameters, and training procedures are also included in the Appendix. The general terminology used in this work is also explained in the Appendix.
@@
Following standard practice in studying memorized models (e.g. Stephenson et al. (2021)), we do not use explicit regularizers such as dropout or batchnorm, or early stopping, unless otherwise mentioned, as a result of which our baseline test accuracy numbers are often much lower than what is usually found with standard training of these models. All the models are trained to either reach very high training accuracies (i.e. $99\%-100\%$) or trained until 500 epochs. Some models did not reach such high accuracies, in which case, results have been shown on the model obtained at epoch 500. We trained 3 instances of each model and results displayed are averaged over these instances with the shaded region indicating the range of results also indicated in the plots.
3.3 Minimum Angle Subspace Classifier Algorithm (MASC)
For a given data point $\bm{x}$ from the training or test set, a layer output data point $\bm{x_{l}}$ from layer $l$ when input $\bm{x}$ is passed through the network and its corresponding training subspaces $\{S_{k}\}_{k=1}^{K}$ with $K$ classes, we use Minimum Angle Subspace Classifier (MASC) Algorithm 1 for predicting class labels $y(\bm{x_{l}})$.
Algorithm 1 Minimum Angle Subspace Classifier (MASC)
Input: Training subspaces $\{S_{k}\}_{k=1}^{K}$, layer output data point $\bm{x_{l}}$ from layer $l$ when input $\bm{x}$ is passed through the network and classes $\{C_{k}\}_{k=1}^{K}$.
Output: MASC prediction class label $y(\bm{x_{l}})$ according to layer $l$ .
1:for each class $C_{k}$ do
[…]
3:  Compute the angle $\theta(\bm{x_{l}},\bm{x_{lk}})$ between $\bm{x_{l}}$ and $\bm{x_{lk}}$
4:end for
5:Assign the label $y(\bm{x_{l}})=C_{k}$ where $k=\arg\min_{k}\theta(\bm{x_{l}},\bm{x_{lk}})$
6:Return: label $y(\bm{x_{l}})$
Given training dataset $\mathcal{D}\{(\bm{x_{i}},y_{i})\}_{i=1}^{m}$, where each $\bm{x_{i}}\in\mathbb{R}^{d}$ and $y_{i}\in\{C_{k}\}_{k=1}^{K}$ are input-label pairs, we estimate training subspaces $\{S_{k}\}_{k=1}^{K}$ for all classes $K$, for a given layer $l$ of the neural network using Algorithm 2 and 3. In practice, $S_{k}$ is represented via its principal components, which form a basis for the subspace.
Algorithm 2 Subspaces Estimator for MASC
Input: Training dataset $\mathcal{D}\{(\bm{x_{i}},y_{i})\}_{i=1}^{m}$, where each $\bm{x_{i}}\in\mathbb{R}^{d}$ and $y_{i}\in\{C_{k}\}_{k=1}^{K}$ are input-label pairs, neural network, and layer $l$.
Output: Subspaces $\{S_{k}\}_{k=1}^{K}$ for classes $K$, for given layer $l$.
[…]
2:for each input pair $(\bm{x_{i}},y_{i})$ in $\mathcal{D}$ do
3:  Pass $\bm{x_{i}}$ through the network layers to obtain the output of layer $l$, denoted as $\bm{x_{l}}\in\mathbb{R}^{ld}$.
4:  $\mathcal{D}_{l}$= $\mathcal{D}_{l}\cup\{\bm{(x_{l},y_{i})}\}$
5:end for
7:Return: Subspaces $\{S_{k}\}_{k=1}^{K}$
Algorithm 3 PCA-Based Subspace Estimation
Input: Layer output $\mathcal{D}_{l}=\{(\bm{x_{i}},y_{i})\}_{i=1}^{m}$, where $\bm{x_{l}}\in\mathbb{R}^{ld}$ and $y_{i}\in\{C_{k}\}_{k=1}^{K}$.
Output: Subspaces $\{S_{k}\}_{k=1}^{K}$ for classes $K$.
[…]
2:for each $\bm{(x_{i},y_{i})}\in\mathcal{D}_{l}$ do
3:  $\mathcal{D}_{\text{new}}\leftarrow\mathcal{D}_{\text{new}}\cup\{\bm{(-x_{i},y_{i}})\}$
4:end for
5:for each $k\in\{1,\ldots,K\}$ do
6:  Extract the subset of data $\mathcal{D}_{\text{new},k}=\{\bm{x_{i}}\mid y_{i}=C_{k}\}$
7:  $S_{k}$ =PCA($\mathcal{D}_{\text{new},k}$)
8:end for
9:Return: Subspaces $\{S_{k}\}_{k=1}^{K}$
We have used 99% as the percentage of variance explained by the principal components, unless otherwise mentioned. While the subspaces are estimated using the training data alone, accuracy of the MASC is determined for the training data and the test data separately. This process is followed for all the layers in the network independently. MASC is using labels of the dataset while creating the class-specific subspaces. The process of creation and use of subspaces with MASC for a new data point are shown schematically in Figure 2.
Figure 2: A schematic of the Minimum Angle Subspace Classifier (MASC) constructed for a specific hidden layer. a) Illustration of the process of fitting a subspace (i.e. a linear space that passes through the origin) corresponding to a single class, for the outputs of a specific hidden layer. For a specific layer, class-wise training dataset (Class 1) is passed through the model till the hidden layer in question. For every output datapoint (activation values) of the layer, a negative data point is added to the point set in order to zero-center the outputs / dataset before performing Principal Components Analysis (PCA).
b) Such subspaces for the hidden layer are constructed for every class. When a new (e.g. unseen) datapoint needs to be classified by MASC, its output from the hidden layer is computed, which is a datapoint in the ambient space of the hidden layer. This datapoint is projected onto the individual class-conditional subspaces and the angle between the data point and its respective projections, $\theta_{1},\ldots\theta_{k}$, are determined. MASC predicts the datapoint’s class to be the one whose subspace the datapoint has the smallest such angle with, i.e $\arg\min_{i}\{\theta_{i}\}$.